The unheated gas in the above system has a volume of 20.0 L at a temperature of 25.0 °C and a pressure of 1.00 atm. The gas is heated to a temperature of 100.0 °C while the volume remains constant. What is the pressure of the heated gas?
1
Expert's answer
2020-05-01T14:30:16-0400

According to the combined gas law equation:

p1V1/T1 = p2V2/T2

where p1 - initial pressure, V1 - initial volume, T1 - initial temperature, p2 - final pressure, V2 - final volume, T2 - final temperature.

If the volume remains constant:

p1/T1 = p2/T2

From here, the pressure of the heated gas is:

p2 = p1T2 / T1

As T1 = 25.0 °C = 298.15 K and T2 = 100.0 °C = 373.15 K:

p2 = 1.00 atm × 373.15 K / 298.15 K = 1.25 atm


Answer: 1.25 atm